OpenStudy (anonymous):

i want to verify sin^4x - cos^4x = sin^2x - cos^2x. HOW DO YOU DO THAT? help me!!!!!

OpenStudy (anonymous):

(sin² x )² - ( cos²x)² = ....

OpenStudy (anonymous):

.....im lost><

OpenStudy (anonymous):

You don't know the formula: a² - b² = ...?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

!!! :'( a² - b² = ( a - b) ( a + b) Here a = sin²x , b = cos²x Now your turn to plug them in!

OpenStudy (anonymous):

(sin² x )² - ( cos²x)² =sin²- cos²x !!!

OpenStudy (anonymous):

Match them up, will you!!! a² - b² = ( a - b) ( a + b) (sin² x )² - ( cos²x)² = sin²- cos²x

OpenStudy (anonymous):

(sin² x )² - ( cos²x)² =(sin²- cos²x) (sin² + cos²x)

OpenStudy (anonymous):

There you go! (sin² + cos²x) = .....?

OpenStudy (anonymous):

(sin² x )² - ( cos²x)²

OpenStudy (anonymous):

You need to do whatever to know this formula, check your text, notes ... sin²x + cos²x =

OpenStudy (anonymous):

1!!!

OpenStudy (anonymous):

Finally (Sighs...)

OpenStudy (anonymous):

Ahhh now i got it! Thank you so muchhh!!! you saved me.

OpenStudy (anonymous):

trig is so hard...thank you!!

OpenStudy (anonymous):

Here, there are 2 BASIC formulas, just 1 of them is trig. and so easily to remember :/